Quantitative Problem: Winston Inc. is trying to determine the effect of its inventory turnover ratio and days sales outstanding on its cash conversion cycle. Winston's 2015 sales (all on credit) were $143,000 and its cost of goods sold was 75% of sales. It turned over its inventory 8.66 times during the year. Its receivables balance at the end of the year was $13,143.85 and its payables balance at the end of the year was $7,398.31. Using this information calculate the firm's cash conversion cycle. Do not intermediate calculations. Round your answer to the nearest whole number.
days
Answer #1

Annual Sales = $143000

Cost of Goods Sold = 0.75*143000 = $107250

Inventory Turnover = 8.66

Accounts Receivable = $13143.85

Accounts Payable = $7398.31

Cash Conversion Cycle = DIO + DSO - DPO

Days Inventory Outstanding (DIO) = 365/Inventory Turnover = 365/8.66 = 42.15 days

Days Sales outstanding (DSO) = (Accounts Receivable/Sales)*365 = (13143.85/143000)*365 = 33.55 days

Days Payables Outstanding (DPO) = (Average Accounts Payables/Cost of Goods Sold)*365 = (7398.31/107250)*365 = 25.18 days

Hence, Cash Conversion Cycle = 42.15 + 33.55 - 25.18 = 50.52 days